# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ C.3 Proof of Theorem 3
$\bigl{(}\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\bigr{)}_{s,t}$ $=\sum_{k_{1}=1}^{n_{i_{0}}}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}+\sum_{k_{2}=1}^{n_{i_{L}}}\mathbf{E}^{({i_{L}})}_{t,k_{2}}\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s}+\mathbf{R}^{(i_{0},i_{L})}_{s,t}+\mathbf{S}^{(i_{0},\dots,i_{L})}_{s,t}.$ (C.8)
@@
$\mathrm{Var}\big{[}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\big{]}=$ $\mathbb{E}\Big{[}\mathrm{Var}\big{[}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\big{|}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}\big{]}\Big{]}+\mathrm{Var}\Big{[}\mathbb{E}\big{[}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\big{|}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}\big{]}\Big{]}$ (C.9)
@@
$\mathrm{Var}\Bigl{[}\sum_{k_{1}=1}^{n_{i_{0}}}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}+\sum_{k_{2}=1}^{n_{i_{L}}}\mathbf{E}^{({i_{L}})}_{t,k_{2}}\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s}\Bigr{]}=$ $\sum_{k_{1}=1}^{n_{i_{0}}}(\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t})^{2}\mathbf{D}^{({i_{0}})}_{s,k_{1}}+\sum_{k_{2}=1}^{n_{i_{L}}}(\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s})^{2}\mathbf{D}^{({i_{L}})}_{t,k_{2}}=\tilde{\sigma}^{2}_{s,t}.$
@@
$\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}:=\tilde{\sigma}^{-1}_{s,t}\mathbf{Y}^{(i_{0})}_{k_{1}}=\tilde{\sigma}^{-1}_{s,t}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}\text{ for any }k_{1}\in[n_{i_{0}}],$
$\tilde{\mathbf{Y}}^{(i_{L})}_{k_{2}}:=\tilde{\sigma}^{-1}_{s,t}\mathbf{Y}^{(i_{L})}_{k_{2}}=\tilde{\sigma}^{-1}_{s,t}\mathbf{E}^{({i_{L}})}_{t,k_{2}}\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s}\text{ for any }k_{2}\in[n_{i_{L}}].$
We now analyze $\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}$ for any fixed $k_{1}\in[n_{i_{0}}]$; the same analysis also applies to $\tilde{\mathbf{Y}}^{(i_{L})}_{k_{2}}$ for any fixed $k_{2}\in[n_{i_{L}}]$. Rewrite $\mathbf{E}^{(i_{0})}_{s,k_{1}}$ as
$\mathbf{E}^{(i_{0})}_{s,k_{1}}=\mathbf{A}^{(i_{0})}_{s,k_{1}}-\mathbf{P}^{(i_{0})}_{s,k_{1}}=[\mathbf{P}^{(i_{0})}_{s,k_{1}}\cdot\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}-\mathbf{P}^{(i_{0})}_{s,k_{1}}]+\mathbf{N}^{(i_{0})}_{s,k_{1}}\cdot\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}.$ (C.10)
Then by Eq. (C.10) we have
$\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}$ $=\underbrace{\tilde{\sigma}^{-1}_{s,t}[\mathbf{P}^{(i_{0})}_{s,k_{1}}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}-\mathbf{P}^{(i_{0})}_{s,k_{1}}]\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}}_{\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}}+\underbrace{\tilde{\sigma}^{-1}_{s,t}[\mathbf{N}^{(i_{0})}_{s,k_{1}}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}]\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}}_{\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}}.$
@@
$\tilde{\sigma}_{s,t}^{-1}\lesssim\zeta_{i_{0},i_{L}}^{-1/2}=\min\Bigl{\{}\frac{(q_{i_{0}}n_{i_{0}}\mu_{i_{0}})^{1/2}}{(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{(i_{0})}\|_{\max}^{2})^{1/2}\|\bm{x}_{i_{L},t}\|},\frac{(q_{i_{L}}n_{i_{L}}\mu_{i_{L}})^{1/2}}{(\sigma_{i_{L}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{(i_{L})}\|_{\max}^{2})^{1/2}\|\bm{x}_{i_{0},s}\|}\Bigr{\}},$ (C.11)
@@
$|\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}|\leq\|\bm{x}_{i_{0},k_{1}}\|\cdot\|\bm{\Lambda}^{(i_{0})}\|^{-1}\cdot\|\bm{x}_{i_{L},t}\|$ $\lesssim(n_{i_{0}}\mu_{i_{0}})^{-1}\|\bm{x}_{i_{0},k_{1}}\|\times\|\bm{x}_{i_{L},t}\|.$ (C.12)
@@
$\mathbb{E}\big{[}|\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|>\epsilon\}\big{]}$ $\leq\mathbb{E}\big{[}|\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|>\epsilon/2\}\big{]}+\mathbb{E}\big{[}|\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\}\big{]}$ (C.13)
$\leq\mathbb{E}\big{[}|\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|>\epsilon/2\}]$
$+\mathbb{E}\big{[}|\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\big{\}}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1]\cdot q_{i_{0}},$
where the last inequality follows from the fact that $\tilde{\mathbf{Y}}^{(i_{0},2)}_{k_{1}}=0$ whenever $\bm{\Omega}_{s,k_{1}}^{(i_{0})}=0$.
Now if $\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1$ then by Eq. (C.11) and Eq. (C.12) we have
$|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|$ $\leq\tilde{\sigma}^{-1}_{s,t}\cdot|\mathbf{P}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}-\mathbf{P}^{(i_{0})}_{s,k_{1}}|\cdot|\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}|\lesssim\frac{(1-q_{i_{0}})\|\mathbf{P}^{(i_{0})}\|_{\max}\cdot\|\bm{x}_{i_{0},k_{1}}\|}{(n_{i_{0}}q_{i_{0}}\mu_{i_{0}})^{1/2}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})^{1/2}}.$ (C.14)
@@
$|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|$ $\leq\tilde{\sigma}^{-1}_{s,t}\cdot|-\mathbf{P}^{(i_{0})}_{s,k_{1}}|\cdot|\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}|\lesssim\frac{q_{i_{0}}^{1/2}\|\mathbf{P}^{(i_{0})}\|_{\max}\cdot\|\bm{x}_{i_{0},k_{1}}\|}{(n_{i_{0}}\mu_{i_{0}})^{1/2}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})^{1/2}}.$ (C.15)
Eq. (C.14) and Eq. (C.15) together imply
$\sup_{k_{1}\in n_{i_{0}}}|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|\leq\epsilon/2$ (C.16)
asymptotically almost surely, provided by the condition in Eq. 3.10. Returning to Eq. (C.13), note that $\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}$ is a deterministic function of $\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}$ and hence
$\mathbb{E}\big{[}|\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ $\leq 2\big{[}|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|^{2}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}\cdot\mathbb{P}\big{[}|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ (C.17)
$+2\mathbb{E}\big{[}|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}.$
We now bound the terms appearing in the above display. First, by Eq. (C.14), we have
$\big{[}|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|^{2}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ $\lesssim\frac{(1-q_{i_{0}})^{2}\|\mathbf{P}^{(i_{0})}\|_{\max}^{2}\cdot\|\bm{x}_{i_{0},k_{1}}\|^{2}}{n_{i_{0}}q_{i_{0}}\mu_{i_{0}}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})}.$ (C.18)
Next, as $\mathbf{N}^{(i_{0})}_{s,k_{1}}$ is sub-Gaussian with $\|\mathbf{N}^{(i_{0})}_{s,k_{1}}\|_{\psi_{2}}\leq\sigma_{i_{0}}$, we have by a similar analysis to Eq. (C.14) that $\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}$ is also sub-Gaussian with
$\|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}\|_{\psi_{2}}\lesssim\frac{\sigma_{i_{0}}\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}}{(n_{i_{0}}q_{i_{0}}\mu_{i_{0}})^{1/2}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})^{1/2}}=:\nu_{i_{0}}.$
@@
$\mathbb{P}\big{[}|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>t|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ $\leq 2\exp\Bigl{(}\frac{-Ct^{2}}{\nu_{i_{0}}^{2}}\Bigr{)}\text{ for any }t>0;$
@@
$\mathbb{E}\big{[}|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ (C.19)
[…]
@@
Combining Eq. (C.13) and Eq. (C.16) through Eq. (C.19), we have
$\sum_{k_{1}=1}^{n_{i_{0}}}\mathbb{E}\big{[}|\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\tilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|>\epsilon\}\big{]}$ $\lesssim n_{i_{0}}\cdot q_{i_{0}}\Bigl{(}\frac{(1-q_{i_{0}})^{2}\|\mathbf{P}^{(i_{0})}\|_{\max}^{2}\cdot\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}^{2}}{q_{i_{0}}n_{i_{0}}\mu_{i_{0}}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})}+\epsilon^{2}+\nu_{i_{0}}^{2}\Bigr{)}\cdot\exp\Bigl{(}-\frac{C\epsilon^{2}}{4\nu_{i_{0}}^{2}}\Bigr{)}$ (C.20)

### v4

@@ D.3 Proof of Theorem 3
$\bigl{(}\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}-\mathbf{P}_{\mathcal{U}_{i_{0}},\mathcal{U}_{i_{L}}}\bigr{)}_{s,t}$ $=\sum_{k_{1}=1}^{n_{i_{0}}}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}+\sum_{k_{2}=1}^{n_{i_{L}}}\mathbf{E}^{({i_{L}})}_{t,k_{2}}\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s}+\mathbf{R}^{(i_{0},i_{L})}_{s,t}+\mathbf{S}^{(i_{0},\dots,i_{L})}_{s,t}$. (D.8)
@@
$\mathrm{Var}\big{[}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\big{]}$= $\mathbb{E}\Big{[}\mathrm{Var}\big{[}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\big{|}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}\big{]}\Big{]}+\mathrm{Var}\Big{[}\mathbb{E}\big{[}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\big{|}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}\big{]}\Big{]}$ (D.9)
@@
$\mathrm{Var}\Bigl{[}\sum_{k_{1}=1}^{n_{i_{0}}}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}+\sum_{k_{2}=1}^{n_{i_{L}}}\mathbf{E}^{({i_{L}})}_{t,k_{2}}\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s}\Bigr{]}$= $\sum_{k_{1}=1}^{n_{i_{0}}}(\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t})^{2}\mathbf{D}^{({i_{0}})}_{s,k_{1}}+\sum_{k_{2}=1}^{n_{i_{L}}}(\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s})^{2}\mathbf{D}^{({i_{L}})}_{t,k_{2}}=\widetilde{\sigma}^{2}_{s,t}.$
@@
$\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}=\widetilde{\sigma}^{-1}_{s,t}\mathbf{Y}^{(i_{0})}_{k_{1}}=\widetilde{\sigma}^{-1}_{s,t}\mathbf{E}^{({i_{0}})}_{s,k_{1}}\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}\text{ for any }k_{1}\in[n_{i_{0}}],$
$\widetilde{\mathbf{Y}}^{(i_{L})}_{k_{2}}=\widetilde{\sigma}^{-1}_{s,t}\mathbf{Y}^{(i_{L})}_{k_{2}}=\widetilde{\sigma}^{-1}_{s,t}\mathbf{E}^{({i_{L}})}_{t,k_{2}}\mathbf{B}^{(i_{L},i_{0})}_{k_{2},s}\text{ for any }k_{2}\in[n_{i_{L}}].$
We now analyze $\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}$ for any fixed $k_{1}\in[n_{i_{0}}]$; the same analysis also applies to $\widetilde{\mathbf{Y}}^{(i_{L})}_{k_{2}}$ for any fixed $k_{2}\in[n_{i_{L}}]$. Rewrite $\mathbf{E}^{(i_{0})}_{s,k_{1}}$ as
$\mathbf{E}^{(i_{0})}_{s,k_{1}}=\mathbf{A}^{(i_{0})}_{s,k_{1}}-\mathbf{P}^{(i_{0})}_{s,k_{1}}=[\mathbf{P}^{(i_{0})}_{s,k_{1}}\cdot\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}-\mathbf{P}^{(i_{0})}_{s,k_{1}}]+\mathbf{N}^{(i_{0})}_{s,k_{1}}\cdot\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}.$ (D.10)
Then by Eq. (D.10) we have
$\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}$ $=\underbrace{\widetilde{\sigma}^{-1}_{s,t}[\mathbf{P}^{(i_{0})}_{s,k_{1}}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}-\mathbf{P}^{(i_{0})}_{s,k_{1}}]\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}}_{\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}}+\underbrace{\widetilde{\sigma}^{-1}_{s,t}[\mathbf{N}^{(i_{0})}_{s,k_{1}}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}]\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}}_{\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}}.$
@@
$\widetilde{\sigma}_{s,t}^{-1}\lesssim\zeta_{i_{0},i_{L}}^{-1/2}=\min\Bigl{\{}\frac{(q_{i_{0}}n_{i_{0}}\mu_{i_{0}})^{1/2}}{(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{(i_{0})}\|_{\max}^{2})^{1/2}\|\bm{x}_{i_{L},t}\|},\frac{(q_{i_{L}}n_{i_{L}}\mu_{i_{L}})^{1/2}}{(\sigma_{i_{L}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{(i_{L})}\|_{\max}^{2})^{1/2}\|\bm{x}_{i_{0},s}\|}\Bigr{\}},$ (D.11)
@@
$|\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}|\leq\|\bm{x}_{i_{0},k_{1}}\|\cdot\|\bm{\Lambda}^{(i_{0})}\|^{-1}\cdot\|\bm{x}_{i_{L},t}\|$ $\lesssim(n_{i_{0}}\mu_{i_{0}})^{-1}\|\bm{x}_{i_{0},k_{1}}\|\times\|\bm{x}_{i_{L},t}\|.$ (D.12)
@@
$\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|>\epsilon\}\big{]}$ $\leq\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|>\epsilon/2\}\big{]}+\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\}\big{]}$ (D.13)
$\leq\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|>\epsilon/2\}]$
$+\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\big{\}}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1]\cdot q_{i_{0}},$
where the last inequality follows from the fact that $\widetilde{\mathbf{Y}}^{(i_{0},2)}_{k_{1}}=0$ whenever $\bm{\Omega}_{s,k_{1}}^{(i_{0})}=0$.
Now if $\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1$ then by Eq. (D.11) and Eq. (D.12) we have
$|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|$ $\leq\widetilde{\sigma}^{-1}_{s,t}\cdot|\mathbf{P}^{(i_{0})}_{s,k_{1}}/q_{i_{0}}-\mathbf{P}^{(i_{0})}_{s,k_{1}}|\cdot|\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}|\lesssim\frac{(1-q_{i_{0}})\|\mathbf{P}^{(i_{0})}\|_{\max}\cdot\|\bm{x}_{i_{0},k_{1}}\|}{(n_{i_{0}}q_{i_{0}}\mu_{i_{0}})^{1/2}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})^{1/2}}.$ (D.14)
@@
$|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|$ $\leq\widetilde{\sigma}^{-1}_{s,t}\cdot|-\mathbf{P}^{(i_{0})}_{s,k_{1}}|\cdot|\mathbf{B}^{(i_{0},i_{L})}_{k_{1},t}|\lesssim\frac{q_{i_{0}}^{1/2}\|\mathbf{P}^{(i_{0})}\|_{\max}\cdot\|\bm{x}_{i_{0},k_{1}}\|}{(n_{i_{0}}\mu_{i_{0}})^{1/2}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})^{1/2}}.$ (D.15)
Eq. (D.14) and Eq. (D.15) together imply
$\sup_{k_{1}\in[n_{i_{0}}]}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|\leq\epsilon/2$ (D.16)
asymptotically almost surely, provided by the condition in Eq. (3.13). Returning to Eq. (D.13), note that $\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}$ is a deterministic function of $\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}$ and hence
$\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ $\leq 2\big{[}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|^{2}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}\cdot\mathbb{P}\big{[}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ (D.17)
$+2\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}.$
We now bound the terms appearing in the above display. First, by Eq. (D.14), we have
$\big{[}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},1)}|^{2}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ $\lesssim\frac{(1-q_{i_{0}})^{2}\|\mathbf{P}^{(i_{0})}\|_{\max}^{2}\cdot\|\bm{x}_{i_{0},k_{1}}\|^{2}}{n_{i_{0}}q_{i_{0}}\mu_{i_{0}}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})}.$ (D.18)
Next, as $\mathbf{N}^{(i_{0})}_{s,k_{1}}$ is sub-Gaussian with $\|\mathbf{N}^{(i_{0})}_{s,k_{1}}\|_{\psi_{2}}\leq\sigma_{i_{0}}$, we have by a similar analysis to Eq. (D.14) that $\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}$ is also sub-Gaussian with
$\|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}\|_{\psi_{2}}\lesssim\frac{\sigma_{i_{0}}\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}}{(n_{i_{0}}q_{i_{0}}\mu_{i_{0}})^{1/2}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})^{1/2}}=:\nu_{i_{0}}.$
@@
$\mathbb{P}\big{[}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>t|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ $\leq 2\exp\Bigl{(}\frac{-Ct^{2}}{\nu_{i_{0}}^{2}}\Bigr{)}\text{ for any }t>0;$
@@
$\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|>\epsilon/2\}|\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}$ (D.19)
$=\int_{t=0}^{\infty}\mathbb{P}\big{[}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|^{2}\cdot\mathbb{I}\big{\{}\|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}\|^{2}>\epsilon^{2}/4\big{\}}\geq t\big{|}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}dt$
$=\epsilon^{2}/4\times\mathbb{P}\big{[}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|^{2}\geq\epsilon^{2}/4\big{|}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}+\int_{t=\epsilon^{2}/4}^{\infty}\mathbb{P}\big{[}|\widetilde{\mathbf{Y}}_{k_{1}}^{(i_{0},2)}|^{2}\geq t\big{|}\mathbf{\Omega}^{(i_{0})}_{s,k_{1}}=1\big{]}dt$
@@
Combining Eq. (D.13) and Eq. (D.16) through Eq. (D.19), we have
$\sum_{k_{1}=1}^{n_{i_{0}}}\mathbb{E}\big{[}|\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|^{2}\cdot\mathbb{I}\{|\widetilde{\mathbf{Y}}^{(i_{0})}_{k_{1}}|>\epsilon\}\big{]}$ $\lesssim n_{i_{0}}\cdot q_{i_{0}}\Bigl{(}\frac{(1-q_{i_{0}})^{2}\|\mathbf{P}^{(i_{0})}\|_{\max}^{2}\cdot\|\mathbf{X}_{\mathcal{U}_{i_{0}}}\|_{2\to\infty}^{2}}{q_{i_{0}}n_{i_{0}}\mu_{i_{0}}(\sigma_{i_{0}}^{2}+(1-q_{i_{0}})\|\mathbf{P}^{i_{0})}\|_{\max}^{2})}+\epsilon^{2}+\nu_{i_{0}}^{2}\Bigr{)}\cdot\exp\Bigl{(}-\frac{C\epsilon^{2}}{4\nu_{i_{0}}^{2}}\Bigr{)}$ (D.20)